OpenStudy (anonymous):

Solve the system 3x+2y=1 and 2x+y=5

OpenStudy (anonymous):

x=9 & y=-13

OpenStudy (amistre64):

solve it by what method? any you perfer?

OpenStudy (anonymous):

any method would be great.... kindly show me how... i don't get it..

OpenStudy (anonymous):

need a walkthrough? you need to eliminate one of x & y from the two equations. i'd suggest you multiplied eqn 2 by 2 & subtract eqn 1 from it. this gives x=9. the rest is about substituting values int the other equation. btw your calculator can do this in less time than it takes to load this page!

OpenStudy (anonymous):

For the conoisseur theres gauss elim, cramers rule & others. for us.. common sense :P

OpenStudy (amistre64):

elimination (simply get rid of a variable) 3x+2y=1 2x+ y=5 <-- multiply by -2 (why?) 3x+2y = 1 -4x-2y =-10 ------------ ; add together -1x = -9 ; and solve for x x = -9/-1 = 9

OpenStudy (anonymous):

or u can use substitution method!!

OpenStudy (amistre64):

sub is a good method; time consuming and labor intensive; but a good method

OpenStudy (anonymous):

substitution comes only after he has eliminated a variable to get the value of the other,

OpenStudy (amistre64):

substituion is the method of solving one of the equations for x or y; then "substituting" that quasi value into the other equation; its elimination of a variable but the long way

OpenStudy (amistre64):

now since x = 9 from above we can take any equation and solve for y with it 2x+y=5 2(9)+y=5 18+y=5 y = -18+5 y = -13

OpenStudy (amistre64):

if you have any questions, just ask :)

OpenStudy (anonymous):

Thank you!!! I get it now!! :D